### v1

# AfriHate: A Multilingual Collection of Hate Speech and Abusive Language Datasets for African Languages

Abstract
Hate speech and abusive language are global phenomena that need socio-cultural background knowledge to be understood, identified, and moderated. However, in many regions of the Global South, there have been several documented occurrences of (1) absence of moderation and (2) censorship due to the reliance on keyword spotting out of context. Further, high-profile individuals have frequently been at the center of the moderation process, while large and targeted hate speech campaigns against minorities have been overlooked. These limitations are mainly due to the lack of high-quality data in the local languages and the failure to include local communities in the collection, annotation, and moderation processes. To address this issue, we present AfriHate: a multilingual collection of hate speech and abusive language datasets in 15 African languages. Each instance in AfriHate is annotated by native speakers familiar with the local culture. We report the challenges related to the construction of the datasets various classification baseline results with and without using LLMs. [^1] The datasets, individual annotations, and hate speech and offensive language lexicons are available on https://github.com/AfriHate/AfriHate
Content Warning: This paper contains representative examples of hate speech and offensive language.
AfriHate: A Multilingual Collection of Hate Speech and Abusive Language Datasets for African Languages
Shamsuddeen Hassan Muhammad1,2thanks: Equal contribution, Idris Abdulmumin3∗, Abinew Ali Ayele4,21, David Ifeoluwa Adelani5, Ibrahim Said Ahmad2,6, Saminu Mohammad Aliyu2, Nelson Odhiambo Onyango7, Lilian D. A. Wanzare7, Samuel Rutunda8, Lukman Jibril Aliyu9, Esubalew Alemneh10, Oumaima Hourrane11, Hagos Tesfahun Gebremichael4, Elyas Abdi Ismail10, Meriem Beloucif12, Ebrahim Chekol Jibril13, Andiswa Bukula14, Rooweither Mabuya14, Salomey Osei15, Abigail Oppong16, Tadesse Destaw Belay17,23, Tadesse Kebede Guge18, Tesfa Tegegne Asfaw4, Chiamaka Ijeoma Chukwuneke19, Paul Röttger20, Seid Muhie Yimam21, Nedjma Ousidhoum22 1Imperial College London, 2Bayero University Kano, 3DSFSI, University of Pretoria, 4Bahir Dar University, 5Mila, McGill University & Canada CIFAR AI Chair, 6Northeastern University, 7Maseno University, 8Digital Umuganda, 9HausaNLP, 10Haramaya University, 11Al Akhawayn University, 12Uppsala University, 13Istanbul Technical University, 14SADiLaR, 15University of Deusto, 16Independent Researcher, 17Instituto Politécnico Nacional, 18Addis Ababa University, 19Lancaster University, 20Bocconi University, 21University of Hamburg, 22Cardiff University, 23Wollo University Contact: s.muhammad@imperial.ac.uk, seid.muhie.yimam@uni-hamburg.de
1 Introduction
No one is born hating another person because of the color of his skin, or his background, or his religion. People must learn to hate, and if they can learn to hate, they can be taught to love, for love comes more naturally to the human heart than its opposite.Mandela (1994)
Table 1: Examples of hateful and abusive instances in AfriHate. All hateful posts are assigned targets.
Hate speech and abusive language are global phenomena that highly depend on specific socio-cultural contexts. Although hate speech and abusive language deviate from the norm on social media, they quickly attract significant attention, spread among online communities Mathew et al. (2019), and incite harm or violence on individuals in real life Saha et al. (2019). Tangible efforts to address these problems must take social and cultural contexts into account Shahid and Vashistha (2023). However, in the absence of high-quality data or when excluding local voices from the collection and annotation processes, one may fail to build assistive tools that help address the problem and moderate such content.
Collecting hate speech and offensive language datasets is complex and time-consuming as researchers typically rely on keywords, hashtags, or user accounts to build datasets Ousidhoum et al. (2020). They may need further insights from both moderators Arora et al. (2023) and affected communities Maronikolakis et al. (2022). Further, resources in languages other than English are scarce, especially for low-resource languages.
To bridge the current gap in the area, we present AfriHate a collection of new hate speech and abusive language datasets in 15 languages spoken in various African regions: Algerian Arabic, Amharic, Igbo, Kinyarwanda, Hausa, Moroccan Arabic, Nigerian Pidgin, Oromo, Somali, Swahili, Tigrinya, Twi, isiXhosa, Yorùbá, and isiZulu. The datasets are annotated by native speakers and include three classes: hate, abusive/offensive, or neutral–neither hateful nor abusive. The targets of the hateful tweets were further labeled based on six common attributes used to discriminate against people: ethnicity, politics, gender, disability, religion, or other. Table 1 shows a sample of the datasets in various languages.
We report the data collection and annotation strategies and challenges when building AfriHate, present various classification baselines with and without using LLMs, and discuss the results. We find that the performance highly depends on the language and that multilingual models can help us boost the performance in low-resource settings. We publicly release the datasets, and individual labels, in addition to manually curated hate speech and offensive language lexicons. These provide a valuable foundation for the research community interested in hate speech and abusive language, African languages, and researchers interested in studying disagreements.
2 Related Work
Language Code Subregion Spoken in Script
Algerian Arabic/Darja arq North Africa Algeria Arabic
Amharic amh East Africa Ethiopia, Eritrea Ethiopic
Hausa hau West Africa Northern Nigeria, Niger, Ghana, and Cameroon, Latin
Igbo ibo West Africa Southeastern Nigeria Latin
Kinyarwanda kin East Africa Rwanda Latin
Moroccan Arabic/Darija ary North Africa Morocco Arabic/Latin
Nigerian Pidgin pcm West Africa Nigeria, Ghana, Cameroon, Latin
Oromo orm East Africa Ethiopia, Kenya, Somalia Latin
Somali som East Africa Somalia, Ethiopia, Djibouti, Kenya Latin
Swahili swa East Africa Kenya, Tanzania, Uganda, DR Congo, Rwanda, Burundi, Mozambique Latin
Tigrinya tir East Africa Ethiopia, Eritrea Ethiopic
Twi twi West Africa Ghana Latin
Xhosa tso Southern Africa Mozambique, South Africa, Zimbabwe, Eswatini Latin
Yorùbá yor West Africa Southwestern and Central Nigeria, Benin, and Togo Latin
Zulu zul Southern Afric Southern Africa Latin
Table 2: Information about the AfriHate languages: their ISO codes, subregions, countries in which they are mainly spoken, and the writing scripts included in AfriHate.
The fast-spreading nature of hate speech and abusive language have been at the center of a significant amount of NLP work in recent years Talat and Hovy (2016); Vigna et al. (2017); Basile et al. (2019); Mansur et al. (2023). However, as there is no unanimous definition of hate speech, researchers in the area have adopted different ones when building resources. For instance, some studies define hate speech as any speech that can cause danger or harm to disadvantaged groups Davidson et al. (2017), others focus on whether the speech is intended to promote hatred Gitari et al. (2015), or whether it dehumanises protected groups Vidgen et al. (2021), which leads to various challenges such as the lack of generalisability Yin and Zubiaga (2021).
Despite Africa being home to more than 2,000 languages and the increasing interest in building hate speech and offensive languages resources for non-English languages Ousidhoum et al. (2019); Madeddu et al. (2023); Röttger et al. (2022), few datasets focus on African languages, such as Amharic Ayele et al. (2024, 2022), Afaan Oromo Ababu and Woldeyohannis (2022), Yorùbá Ilevbare et al. (2024), Hausa Vargas et al. (2024); Adam et al. (2023), and Nigerian Pidgin Ndabula et al. (2023); Ilevbare et al. (2024); Aliyu et al. (2022); Tonneau et al. (2024). Moreover, most resources adopt a binary labeling scheme (hate/offensive) and do not include a normal class (e.g., Aliyu et al. (2022)), or do not add the target attributes (e.g., Ilevbare et al. (2024)). Other work Tonneau et al. (2024) relies on active learning for annotating some data instances, which is not ideal when labeling hate speech for under-resourced African languages or when focusing on underrepresented cultures Lee et al. (2023, 2024). Similarly, a limited number of studies involve African communities in the dataset creation process (Adelani et al., 2021; Maronikolakis et al., 2022; Abdulmumin et al., 2024). We take a step forward in addressing this problem by developing 15 new datasets for hate and offensive speech in various languages spoken across the African continent.
amh arq ary hau ibo kin oro pcm som swa tir twi xho yor zul
Hate 88 62 41 36 46 264 126 23 45 24 58 20 42 68 31
Abusive 74 40 0 149 118 362 159 26 67 12 66 86 177 109 118
Table 3: Number of keywords used for data collection. For Algerian Arabic (arq), the collection includes controversial topics that are neither hateful nor offensive.
3 Creating AfriHate
AfriHate covers 15 languages from various African regions. In Table 2, we report the scripts of these languages and the main regions where they are spoken. The collection includes tweets from 2012 to 2023 collected using the Academic API before the suspension of free academic access. The API allowed us to collect up to 10 million tweets per month, and Twitter/X is a commonly used platform in African countries with documented cases of hate speech propagation Adjai and Lazaridis (2013); Egbunike et al. (2015); Oriola and Kotzé (2020); Ridwanullah et al. (2024); Raborife et al. (2024).
3.1 Data Collection
Except for Amharic and Tigrinya, the Twitter API does not support African languages, which makes the data collection challenging. We, therefore, follow strategies adopted by previous work such as Muhammad et al. (2022, 2023) and use various heuristics based on hate speech and abusive language keywords, user handles, stopwords, hashtags, and locations. Table 3 shows the number of keywords used for data collection in each language.
Since interpreting hate and abusive content heavily depends on understanding political and socio-cultural contexts, we have adopted language-dependent collection and annotation strategies. As previous work relied on specific keywords to collect data (e.g., Talat and Hovy, 2016; Basile et al., 2019), we follow a similar strategy by using a larger set of keywords. Similarly to Ousidhoum et al. (2019), our keywords include culture-specific controversial topics and are of various sizes (see Table 3). However, for some languages such as Nigerian-Pidgin and Hausa, an initial pre-annotation phase revealed a limited number of hateful tweets. We therefore used additional strategies to collect more tweets: 1) keyword crowdsourcing, 2) manual data collection, and 3) using existing datasets, as we explain in the following.
Crowdsourcing Keywords
To crowdsource additional keywords, we first asked native speakers to provide us with a list of hateful, abusive, or controversial keywords. To further diversify our lists, we contacted social media influencers, who asked their followers to share abusive or hateful keywords in their local languages by filling in a form created for anonymous collection. This strategy resulted in a broader vocabulary as the followers come from various backgrounds. Moreover, we used off-the-shelf curated hate speech lexicons from the PeaceTech Lab [^2] https://www.peacetechlab.org/the-peacetech-toolbox and the Hatebase. [^3] https://hatebase.org The lists were post-processed by native speakers prior to the collection of the tweets.
Manual Data Collection
For Kinyarwanda and Twi, native speakers manually collected all the tweets using a combination of keywords and user handles. We curated a list of user handles of public figures who frequently post hateful or abusive content, and collected tweets from their profiles.
Using Existing Datasets
For Nigerian Pidgin, we used an additional existing hate speech dataset from Tonneau et al. (2024). Since some instances in this dataset were annotated using active learning, we re-annotated those labeled as hateful or offensive only. Similarly, for Swahili, we selected instances from the sentiment analysis dataset by Muhammad et al. (2023) [^4] We selected only negative instances, as they may contain hateful or abusive content., the misinformation dataset by Amol et al. (2023), and the hate speech one by Ombui et al. (2019).
These instances were re-annotated into our pre-define classes (hate, offensive and normal). We further labeled the targets of the tweets based on our pre-defined target attributes, i.e., disability, ethnicity, gender, politics, religion, and others.
3.2 Data Processing
We further cleaned the collected tweets and removed retweets, tweets containing less than three words, duplicates, URLs, invisible characters, and redundant white spaces. We converted the tweets written in Latin script to lowercase and anonymized the tweets by replacing @mentions with a placeholder @user.
3.3 Language Identification
We collected the tweets using location and keywords. This is particularly challenging for African languages since people within one location can speak different languages. That is, keywords and hashtags may appear in more than one language, which makes the data selection more difficults
Although ptudies have use Muhammad et al. (2022, 2023) open-source and closed-source language identification (LID) tools, these often show low accuracy in African languages, especially when used in social media posts. This is largely due to the unique linguistic characteristics of these languages, such as the common usage of code-mixing and digraphia, i.e., a language can be written in more than one script.
To address these limitations, we built a LID model that improves identification performance for social media text data in our target languages through continued pretraining of AfroXLMR model Alabi et al. (2022) on Glot500-c corpus which covers 511 predominantly low-resource languages Imani et al. (2023). We further fine-tuned the model on AfriSenti-LID dataset, which focuses on social media data and covers most of the languages included in AfriHate and a similar Twitter-sphere. Our LID tool and its documentation can be found on our project page. [^5] https://github.com/hausanlp/AfriLID
Language amh arq ary hau ibo kin oro pcm som swa tir twi xho yor zul
Manually Collected
Pre-Annotation
Total Annotators 11 7 3 3 6 3 9 3 7 5 8 3 3 4 3
Annotators per Instance 4 3 3 3 3 3 4 3 4 3 4 3 3 4 3
Free-Marginal Multirater Kappa $\uparrow$ 0.63 0.68 0.61 0.75 0.80 0.81 0.63 0.65 0.46 0.55 0.46 0.75 0.62 0.68 0.81
Table 4: Collection and annotation details for the AfriHate dataset. The table shows if the data was manually collected, whether a pre-annotation step was conducted, the total number of annotators, the number of annotators per instance, and the inter-annotator agreement (Free-Marginal Multirater Kappa).
Class amh ary arq hau ibo kin oro pcm som swa tir twi xho yor zul
Hate 2,246 162 674 343 251 1,268 2,293 1,177 388 3,974 2,940 443 210 150 147
Abusive 1,353 778 2,270 2,336 3,510 1,146 667 5,238 1,404 7,708 1,070 3,278 1,550 2,655 1,839
Neutral 1,359 310 1,690 3,965 1,242 2,308 2,072 4,184 2,868 9,410 1,062 180 1,923 2,074 2,322
Table 5: Number of instances per class in each dataset.
3.4 Data Annotation
3.4.1 Pre-Annotation and Data Selection
We randomly sampled tweets in each language and conducted a pilot annotation, which showed a large class imbalance despite collecting tweets using abusive and hateful keywords. For instance, most tweets in Hausa were neutral (neither hateful nor abusive) due to keywords carrying multiple meanings depending on the region where the word is used, i.e., it can have a neutral connotation in some parts of Nigeria. For example, the word Aboki means “friend” in Northern Nigeria, while it can be an insult in the Southern part of the country. As this would have led to an insufficient number of instances in each target class, we included a pre-annotation phase to ensure each class covered a reasonable percentage of the data.
During the pre-annotation, we provided the annotators with a distinct pool of tweets and asked them to select those likely to be hateful or abusive. We then aggregated the pre-selected tweets, and multiple annotators labeled them. Table 4 indicates the languages for which we conducted a pre-annotation step, i.e., only those for which we observed a significantly high imbalance during the pilot annotation.
Hate Target amh arq ary hau ibo kin oro pcm som swa tir twi xho yor zul
Disability 0 0 26 0 0 6 0 1 0 112 0 32 3 2 0
Ethnicity 902 269 150 32 383 94 462 537 74 2,799 573 245 94 103 241
Gender 15 2 6 2 3 24 12 50 17 117 0 12 111 13 0
Politics 1,247 118 63 0 6 1,096 1,550 87 703 233 2,251 32 0 33 0
Religion 130 6 207 38 0 10 31 105 19 336 16 14 0 22 0
Others 199 0 0 0 0 0 767 15 208 501 85 14 0 76 12
Table 6: Data distribution of hate speech targets in AfriHate.
Split amh ary arq hau ibo kin oro pcm som swa tir twi xho yor zul
Train 3,467 3,240 716 4,566 3,419 3,302 3,517 7,416 3,174 14,760 3,547 2,564 2,502 3,336 2,940
Dev 744 695 211 1,029 774 706 763 1,590 741 3,164 760 639 559 724 640
Test 747 699 323 1,049 821 714 759 1,593 745 3,168 765 698 622 819 728
Table 7: Number of instances included in the training (train), development (dev), and test splits of the different datasets.
3.4.2 Recruiting Annotators
The unavailability of annotators for African languages on common platforms like Amazon Mechanical Turk and Prolific makes traditional crowdsourcing methods impractical. As Kirk et al. (2023) demonstrated that trained annotators achieve higher quality results, we trained native speakers and recruited them as annotators. For each language, we also recruited a native speaker as a language lead who would control for the quality of the annotations. We used Label Studio as an annotation platform [^6] https://labelstud.io/ and an adapted version of the Potato annotation tool [^7] https://github.com/davidjurgens/potato.
3.4.3 Annotation Task
Labels
We provided the annotators with thorough guidelines (see Figure 1 in the Appendix). We asked the annotators to choose one of three categories: Hate, Abusive/Offensive, or Neutral. The latter means that the tweet is neither hateful nor abusive. Tweets spotted in a language different from the target one were labeled Indeterminate and were later excluded from the final dataset.
Similarly to Ayele et al. (2024); Ousidhoum et al. (2019); Fortuna et al. (2019), annotators had to select the target(s) of the hateful tweets. That is, the common attribute(s) based on which the tweet is discriminating against people: Ethnicity, Politics, Gender, Disability, Religion, or Other. We do not include targets offensive and abusive tweets as these can often be generic and directed towards an individual as previously reported by Zampieri et al. (2019) (e.g., see the Algerian Arabic example in Table 1). Details about the targets can be found in Figure 1 in the Appendix. For some languages such as Hausa, we asked the annotators to spot which words made them label the tweet hateful or abusive.
Final label selection
As shown in Table 4, for languages where we conducted a pre-annotation step, each tweet was annotated by 3 annotators, leading to a total of 4 labels per tweet with the pre-annotation label counting as one. On the other hand, for instances in datasets for which we did not carry a pre-annotation step, 3 to 4 annotators were assigned to each tweet, and the final gold label was determined by majority voting, i.e., two out of three labels or three out of four labels.
Table 5 and Table 6 show the final number of instances per class and the target distributions for all the languages.
3.4.4 Inter-Annotator Agreement
To assess the inter-annotator agreement (IAA), we computed the free marginal Randolph’s Kappa score Randolph (2005) for each of the datasets. We chose this metric as it is suitable for tasks involving multiple annotators.
Table 4 shows the IAA scores for all the datasets, the total number of annotators in each, and the number of annotators per instance. The IAA scores range from $0.46$ to $0.81$, indicating medium to high agreement levels. The highest agreement scores are reported for kin, and twi, which can be due to the manual collection of only potentially abusive and hateful tweets. For other languages such as hau and zul, the high agreement can be attributed to the pre-annotation step, which helped us filter tweets that were later annotated
3.5 Dataset Statistics
As shown in Table 5, most datasets are imbalanced, and the hate class includes fewer instances in 9 out of 15 languages. The variations in the class distributions are due to the differences between the languages and the data collection techniques. Further, the target distributions also differ because of socio-cultural characteristics related to local politics, social dynamics, and an unavoidable degree of selection bias Ousidhoum et al. (2020).
Data Splits
We split the AfriHate datasets based on the various label distributions. As reported in Table Table 7, each test set includes a minimum of 100 instances in each class (i.e., hate, abusive, and normal). This guarantees a more robust evaluation of the different models.
Model amh ary arq hau ibo kin oro pcm som swa tir twi xho yor zul avg.
Monolingual
AfriBERTa 69.54 67.93 30.48 82.28 89.53 79.43 73.43 66.90 65.52 91.36 73.07 74.54 81.07 72.37 83.75 72.33
AfriTeVa V2 73.91 76.71 25.25 79.06 83.95 77.60 71.61 68.69 69.65 90.68 72.36 64.96 54.67 79.88 69.05 68.73
AfroXLMR 70.65 80.16 61.18 81.93 89.30 80.72 72.11 67.98 66.84 91.44 74.52 77.17 82.49 72.15 83.44 76.15
AfroXLMR-76L 74.36 80.05 53.52 82.78 89.59 79.58 76.63 68.38 71.09 91.72 76.27 76.65 84.40 72.35 84.65 76.45
Multilingual
AfroXLMR-76L 75.25 80.76 63.31 82.20 89.85 79.56 77.62 69.20 72.26 91.22 77.55 78.68 86.83 74.32 86.81 78.16
Table 8: Model performances after fine-tuning BERT-based LMs. The best performance for each language is highlighted in bold.
4 Experiments
4.1 Setup
We compare the performance of three main sets of approaches on the AfriHate datasets:
  1. 1.
    Fine-tuning BERT-based pre-trained language model (PLM),
  2. 2.
    Few-shot learning with SetFit with BERT-based PLM (Tunstall et al., 2022): a few-shot approach using BERT-like PLMs,
  3. 3.
    Prompting large language models (LLMs) in zero and few-shot settings.
4.1.1 Fine-tuning PLMs
We use four widely adopted Africa-centric PLMs that have been shown to consistently perform better than massively multilingual PLMs such as XLM-R (Conneau et al., 2020). The models are AfriBERTa-large (Ogueji et al., 2021), AfriTeVa V2 base (Oladipo et al., 2023), AfroXLMR (Alabi et al., 2022) and AfroXLMR-76L (Adelani et al., 2024). Each model was trained for 20 epochs over 5 runs with a batch size of 32, maximum sequence length of 128, and a learning rate of $55e-$ 5 except for AfroXLMR-76L where we used a learning rate of $53e-$ 5. The rest of the hyperparameters were the default values set in the HuggingFace fine-tuning pipeline for text classification.
4.1.2 SetFit Few-shot Learning
SetFit is a few-shot learning approach based on sentence-transformer models like LabSE Feng et al. (2022). It works by, first, fine-tuning a pre-trained sentence transformer model on a few examples in a contrastive manner. Then, the resulting model is used to generate rich text embeddings, which are used to train a classification head. We used LaBSE to train classifiers with the following configurations:
  1. 1.
    for zero-shot learning, we trained the transformers for one epoch using the dummy dataset generated by the framework (2 x [This sentence is {Label}, …]), where {Label} can be “Neutral”, “Abusive” or “Hate”;
  2. 2.
    for few-shot learning, we trained each model for three epochs using 5, 10, and 20 shots. All the classifiers were trained using a batch size of 32.
4.1.3 Prompting LLMs
We prompt one closed model (GPT-4o) and nine open models of various model sizes (0.4B to 70B). The open models are: InkubaLM-0.4B (Tonja et al., 2024), mT0-small (Muennighoff et al., 2023), BLOOMZ 7B  (Scao et al., 2022), Mistral 7B (Jiang et al., 2023), Aya-23-35B (Aryabumi et al., 2024), LLaMa 3.1 {8B & 70B} (Dubey et al., 2024), and Gemma 2 {9B & 27B} (Team et al., 2024). All the models were prompted using five prompt templates with clear definitions of the Abusive, Hate, and Neutral categories. For hate speech, we used the definition adopted by the United Nations and another one from the Merriam Webster’s dictionary. We report the average scores across all the templates in the results Section. The full prompts can be found in Section A.1.
Lang. Monolingual AfroXLMR-76L Multilingual AfroXLMR-76L GPT-4o
Macro F1 Abuse Hate Neutral Macro F1 Abuse Hate Neutral Macro F1 Abuse Hate Neutral
amh 73.83 70.20 77.27 74.02 75.55 71.92 78.18 76.54 65.70 59.21 67.91 70.00
ary 78.01 86.81 66.98 80.25 79.05 86.59 68.66 81.91 75.93 84.29 61.09 82.43
arq 57.05 77.35 27.42 66.38 68.99 78.29 55.95 72.73 73.67 77.63 71.43 71.96
hau 81.53 77.60 80.45 86.54 78.05 76.45 72.83 84.88 59.44 70.61 40.77 66.95
ibo 88.00 92.18 91.18 80.64 88.33 92.41 91.18 81.40 76.85 84.13 75.37 71.04
kin 77.96 70.09 80.87 82.90 77.11 70.00 78.90 82.43 74.27 65.45 82.01 75.37
orm 70.07 47.19 81.17 81.86 69.93 48.89 80.11 80.78 65.30 44.12 72.78 79.00
pcm 65.40 71.09 52.82 72.31 63.71 69.66 49.10 72.39 63.53 57.38 56.67 76.54
som 59.53 67.00 30.77 80.83 60.91 68.81 32.94 81.00 62.94 63.85 44.91 80.05
swa 88.00 92.18 91.18 80.64 89.50 89.35 88.33 90.83 83.29 85.48 80.49 83.91
tir 72.32 72.29 82.60 62.07 74.45 75.23 84.97 63.16 56.23 51.73 64.38 52.59
twi 58.30 91.45 61.40 22.05 63.66 91.04 60.18 39.76 62.64 85.90 54.45 47.58
xho 79.37 86.33 66.23 85.53 81.57 85.14 71.95 87.63 57.74 70.82 38.46 63.94
yor 57.36 84.24 6.90 80.95 59.41 84.03 11.76 82.44 71.48 81.56 53.01 79.88
zul 80.97 87.54 68.79 86.57 84.34 85.46 79.55 88.01 70.63 74.60 67.78 69.51
Table 9: Macro F1 scores per class for monolingual and multilingual AfroXLMR-76L [only one run] vs. GPT-4o [prompt template 1; 20 shots]. The best performance for each language is highlighted in bold.
4.2 Experimental Results
Model # Shots amh ary arq hau ibo kin oro pcm som swa tir twi xho yor zul Avg.
SetFit 0 33.79 46.54 27.4 23.57 36.90 25.29 26.26 43.46 32.18 44.38 49.30 51.13 37.38 48.46 35.08 37.41
5 44.89 33.82 36.07 49.93 51.52 55.35 36.56 53.28 48.09 54.90 35.36 33.23 47.49 34.36 33.18 43.20
10 49.42 35.81 46.55 49.72 39.31 53.10 39.10 57.00 42.29 62.64 39.16 45.76 43.52 49.73 40.33 46.23
20 50.13 40.49 54.24 55.40 65.15 57.35 40.91 59.09 48.95 74.93 40.04 53.72 46.92 56.54 50.67 52.97
Mistral-7B-v0.1 0 21.18 17.46 11.98 6.84 8.04 15.18 20.82 6.76 8.16 18.88 25.62 8.38 9.72 8.22 8.92 13.08
5 37.68 31.60 44.54 34.90 36.84 46.98 39.46 50.00 34.12 58.58 35.64 32.04 34.24 41.04 32.54 39.35
10 39.68 36.50 50.56 37.82 36.94 45.72 39.02 52.10 31.82 63.18 35.84 30.82 32.64 42.90 32.46 40.53
20 36.64 35.46 52.94 38.44 38.98 52.28 39.24 54.60 33.28 66.28 37.42 30.44 34.12 46.40 35.00 42.10
aya-23-35B 0 17.04 19.20 17.20 20.34 14.04 21.34 20.38 20.02 18.56 34.04 17.10 10.68 20.32 17.48 20.24 19.20
5 37.78 56.40 57.00 39.70 42.70 48.78 39.24 57.90 38.72 69.84 35.68 40.24 36.76 48.88 44.58 46.28
10 38.50 53.96 61.82 44.42 45.80 52.82 42.16 59.26 37.60 74.10 35.88 39.92 36.18 48.44 45.00 47.72
20 38.10 52.90 63.76 46.14 46.52 57.62 42.26 61.68 38.68 75.68 38.12 38.54 37.74 51.52 44.52 48.92
Gemma-2-9B 0 27.42 28.36 25.46 14.48 15.00 24.08 23.74 24.74 10.92 37.12 25.52 19.50 13.40 21.66 12.52 21.59
5 56.60 57.68 61.14 49.98 48.12 60.26 45.10 60.30 46.42 74.78 48.40 46.50 35.82 55.96 46.60 52.91
10 57.84 61.08 61.72 56.62 53.78 63.54 45.78 60.54 51.72 78.16 45.48 48.16 38.82 60.22 52.58 55.74
20 60.96 59.94 64.38 55.90 54.56 64.14 44.70 62.54 52.22 79.24 44.94 48.92 39.02 60.50 53.68 56.38
Gemma-2-27B 0 41.78 41.24 41.12 28.96 31.36 42.08 33.46 49.78 31.10 59.88 30.84 27.74 25.64 41.46 28.02 36.96
5 59.62 64.14 65.62 54.62 56.14 61.08 46.76 61.78 54.12 81.18 52.12 47.26 40.48 61.72 54.28 57.39
10 60.70 61.36 64.88 58.90 56.84 64.86 46.96 61.60 52.86 81.94 49.82 50.86 41.90 59.66 56.94 58.01
20 62.28 59.86 65.80 59.60 58.76 66.06 48.90 63.24 52.90 82.98 53.36 50.34 43.08 59.10 56.88 58.88
Llama-3.1-70B 0 36.34 43.34 42.64 35.64 32.52 38.52 31.54 48.66 31.14 60.14 27.08 25.54 28.98 40.64 35.50 37.21
5 58.52 66.24 62.88 52.86 52.88 58.14 45.50 64.38 52.72 75.76 44.42 43.90 39.02 58.08 55.24 55.37
10 61.18 64.46 62.20 56.40 55.10 59.00 46.60 63.58 54.40 78.74 49.62 45.76 39.04 57.96 55.74 56.65
20 60.38 61.36 63.80 57.40 53.80 62.48 49.02 63.18 51.82 80.02 53.90 47.50 40.34 57.94 56.36 57.29
GPT-4o 0 61.78 66.41 73.75 56.91 68.82 62.53 60.01 65.94 63.42 73.66 45.75 52.72 58.92 75.21 54.47 62.69
5 66.73 73.53 77.33 55.44 76.73 72.27 70.94 66.29 59.33 80.95 61.11 73.21 60.45 76.98 59.34 68.71
10 67.94 75.54 77.54 58.15 80.08 75.07 72.27 67.14 62.75 84.19 57.52 72.28 65.75 76.74 65.05 70.53
20 68.08 76.16 78.69 58.34 80.81 74.86 72.33 65.41 66.44 84.61 59.55 75.86 66.08 77.11 71.36 71.71
Lang. avg. - 48.32 50.74 54.04 44.91 47.79 52.88 43.18 54.44 43.10 67.53 41.95 42.53 39.06 51.25 44.18 48.39
Table 10: Model performance (Macro F1-score) for zero- and-few shot classifiers across the different languages in AfriHate. The best performance for each language is highlighted in bold. This is an average over 5 prompt templates.
4.2.1 Monolingual vs. Multilingual Fine-tuning
We compare monolingual fine-tuning where we train on a language and evaluate on the same language to multilingual fine-tuning—where we combine the training data of all the languages and evaluate the results for each.
Table 8 shows the results of the fine-tuning experiments. We find that encoder-only models perform better than the T5-style models, i.e., AfriTeVa V2. On average, AfroXLMR-76L achieved the best performance, most likely due to the fact that it was pre-trained on all the languages included in AfriHate. While AfroXLMR was not pre-trained on some languages such as Tigrinya (tir) and Twi (twi), it still achieves performance that is comparable to AfroXLMR-76L. AfriBERTa generally struggles with Arabic dialects such as Algerian Darja (arq) and Moroccan Darija (ary) as the Arabic script was not included in its pre-training.
Overall, multilingual training of a single model leads to the best results on 11 out of 14 languages, and comparable results on the remaining languages except for Yorùbá (yor), where AfriBERTa led to the best result because of its Africa-centric tokenizer.
Table 9 shows the per-class accuracy across different languages. Overall, multilingual models perform better for languages with a low percentage of the “hate” category in the training data (e.g., $<200$) such as ary, xho, yor, and zul with an F-score improvement of $+1.7$, $+5.7$, $+4.9$, and $+10.8$, respectively.
4.2.2 Zero-shot vs. Few-shot Settings
Table 10 shows the results of both zero-shot and few-shot experiments. SetFit performs slightly better than all open LLMs in zero-shot settings ($36.9$), and GPT-4o leads to the best overall performance with $61.9$ F1 points.
When considering the few-shot settings, 5-shot models show the biggest boost in performance, where Gemma 2 9B and other bigger models (Gemma 2 27B and LLaMa 3.1 70B) improve by about $+20$ points. The performance boost with additional 10 and 20 shots is more limited for LLMs ($+3.0$ improvement), whereas SetFit consistently benefits from additional examples. The best results reached for closed models are at 20-shots, where GPT-4o achieved an overall F1-score of $70.8$ while Gemma 2 27B achieved the best overall results for any open model with an F1-score of $57.2$.
In our performance analysis per different classes shown in Table 9, GPT-4o generally performs worse than full fine-tuning in monolingual or multilingual settings. However, we find that it achieves significantly better performance for hate detection in a few languages such as arq ($+44.0$), som ($+14.9$), and yor ($+46.1$) compared to monolingual fine-tuning. For languages without enough training data for the hate category such as Yorùbá, prompting LLMs might provide a better detection of this class compared to fine-tuning BERT-like PLMs.
4.2.3 Overall Results
Overall, our results show that fine-tuning multilingual models leads to a better performance for the majority of the AfriHate languages. That is, AfroXLMR-76L achieves an average macro F1 score of 78.16. As for zero-shot and few-shot settings, GPT-4o outperformed other models, with average F1 scores of 61.89 and 70.79 in zero-shot and 20-shot settings, respectively.
Overall, these results highlight the advantages of multilingual and context-specific models in hate and abusive language detection for African languages.
5 Conclusion
We introduced AfriHate, the first large-scale collection of hate and abusive language datasets in 15 African languages: Algerian Arabic, Amharic, Igbo, Kinyarwanda, Hausa, Moroccan Arabic, Nigerian Pidgin, Oromo, Somali, Swahili, Tigrinya, Twi, Xhoza, Yorùbá, and Zulu. The datasets were annotated by native speakers as hate speech, abusive, or neutral. We discussed our data collection strategies and highlighted the challenges faced during the data collection and annotation. We then reported baseline experiments using Africa-centric pre-trained language models as well as prompted open and closed LLMs showing a large gap in the performance across languages.
AfriHate is a first step towards building high-quality hate speech resources for African languages. We publicly release all the datasets, scripts, models, and lexicons to the research community.
6 Limitations
While we collected AfriHate using large sets of keywords, we acknowledge the unavoidable presence of selection bias Ousidhoum et al. (2020) as no dataset can capture the full range of hate speech contexts across various languages and cultures. In addition, although we recruited annotators who come from different socio-cultural backgrounds, hate speech remains a subjective task and one cannot include all possible perspectives of what constitutes hate speech or abuse. We mitigate the problem by sharing the individual annotations with the research community studying the problem.
Further, when using language identification to collect data, challenges due to code-mixing and digraphia, make the task non-trivial given the common usage of multilinguality in African languages. We address the problem by asking the annotators to flag any tweet that is not in the target language. We acknowledge, nevertheless, instances that may have been missed by our annotators.
Finally, we report on the various dataset statistics and model features. However, given the fact that we use some closed models in our experiments, and the class imbalance problem which is inherent to hate speech datasets, we do not claim that our results are fully replicable or generalisable.
7 Ethical Considerations
Annotators
The annotators involved in this study were compensated for their work by more than the minimum wage and any demographic information about them was shared with consent. We acknowledge the difficulty of annotating hate speech and abusive language on people’s well-being. Therefore, the annotators could reach out to us and were allowed to quit at any time.
Language Use
Our datasets focus on hate speech and abusive languages in 15 African languages. However, we do not claim that our datasets represent the full usage of these languages. We further acknowledge the socio-cultural biases that can come with the data as views on hate highly differ from one person to another and those shared by our annotators cannot include all possible perspectives.
Intended uses and potential misuses
Our datasets focus on hate speech and abusive language. They present a first step towards studying the phenomenon in some low-resource African languages. However, as malicious data actors can misuse our resources, we follow the suggestions made by Schlichtkrull et al. (2023) for automated fact-checking researchers and clearly state the following:
  • Models built using our datasets should not be used for automated removal.
  • Our data subjects are social media users.
  • Our data actors and model owners should be users, moderators, experts, and researchers with background knowledge in the field, especially on the limitations of automated hate speech and abusive language detection models.
  • Given the sensitivity of the task and the high risk of false positives, any constructed or deployed model using our data should be human-in-the-loop with the humans being native or near-native speakers.
Acknowledgments
We are grateful for the work and insights shared by our annotators.
This work was carried out with support from Lacuna Fund, an initiative co-founded by The Rockefeller Foundation, Google.org, and Canada’s International Development Research Centre. The views expressed herein do not necessarily represent those of Lacuna Fund, its Steering Committee, its funders, or Meridian Institute. We thank Label Studio for providing the academic version annotation tool used in this project. We also thank OpenAI for providing API credits to Masakhane. Shamsuddeen acknowledges the support of Google DeepMind. Paul was supported by a MUR FARE 2020 initiative under grant agreement Prot. R20YSMBZ8S (INDOMITA) and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (No. 949944, INTEGRATOR). Abinew acknowledges the support of LT group, University of Hamburg.
Appendix A Example Appendix
A.1 Prompt templates
Prompt 1 I am providing you with the definition of Hate speech, Abusive language, and Neutral tweets. Hate speech is language content that expresses hatred towards a particular group or individual based on their political affiliation, race, ethnicity, religion, gender, sexual orientation, or other characteristics. It also includes threats of violence. Neutral language does not contain any bad language. Which category does the tweet above belong to: ’Hate’, ’Abuse’, or ’Neutral’? Pick exactly one category. Don’t give any additional context, just classify the tweet. Tweet: text Category:
Prompt 2 Read the following label definitions and provide a label without any explanations. Hate: Hate speech is public speech that expresses hate or encourages violence towards a person or group based on something such as race, religion, gender, ethnicity, sexual orientation or other characteristics. Abusive: Abusive and offensive language means verbal messages that use words in an inappropriate way and may include but is not limited to swearing, name-calling, or profanity. Offensive language may upset or embarrass people because it is rude or insulting Neutral: Neutral language is neither hateful nor abusive or offensive. It does not contain any bad language. Text: tweet Label:
Prompt 3 Read the following text and definitions: Text: tweet. Definitions: Hate: Hate speech is public speech that expresses hate or encourages violence towards a person or group based on something such as race, religion, gender, ethnicity, sexual orientation or other characteristics. Abuse: Abusive and offensive language means verbal messages that use words in an inappropriate way and may include but is not limited to swearing, name-calling, or profanity. Offensive language may upset or embarrass people because it is rude or insulting Neutral: Neutral language is neither hateful nor abusive or offensive. It does not contain any bad language. Which of these definitions (hate, abuse, neutral) apply to this tweet?
Prompt 4 Read the following definitions and text to categorize: Definitions: Hate: Hate speech is public speech that expresses hate or encourages violence towards a person or group based on something such as race, religion, gender, ethnicity, sexual orientation or other characteristics. Abuse: Abusive and offensive language means verbal messages that use words in an inappropriate way and may include but is not limited to swearing, name-calling, or profanity. Offensive language may upset or embarrass people because it is rude or insulting Neutral: Neutral language is neither hateful nor abusive or offensive. It does not contain any bad language. Text: tweet. Which of these definitions (hate, abuse, neutral) apply to this tweet?
Prompt 5 You will be given a text snippet and 3 category definitions. Your task is to choose which category applies to this text. Your text snippet is: tweet Your category definitions are: HATE category definition: Hate speech is public speech that expresses hate or encourages violence towards a person or group based on something such as race, religion, gender, ethnicity, sexual orientation or other characteristics. ABUSE category definition: Abusive and offensive language means verbal messages that use words in an inappropriate way and may include but is not limited to swearing, name-calling, or profanity. Offensive language may upset or embarrass people because it is rude or insulting NEUTRAL category definition: Neutral language is neither hateful nor abusive or offensive. It does not contain any bad language. Does the text snippet belong to the HATE, ABUSIVE, or the NEUTRAL category? Thinking step by step answer HATE, ABUSIVE, or NEUTRAL capitalizing all the letters. Explain your reasoning FIRST, then output HATE, ABUSIVE, or NEUTRAL.
Model # Shots amh ary arq hau ibo kin oro pcm som swa tir twi xho yor zul Avg.
SetFit 0 33.79 46.54 27.40 23.57 36.90 25.29 26.26 43.46 32.18 44.38 49.30 51.13 37.38 48.46 35.08 37.41
5 44.89 33.82 36.07 49.93 51.52 55.35 36.56 53.28 48.09 54.90 35.36 33.23 47.49 34.36 33.18 43.20
10 49.42 35.81 46.55 49.72 39.31 53.10 39.10 57.00 42.29 62.64 39.16 45.76 43.52 49.73 40.33 46.23
20 50.13 40.49 54.24 55.40 65.15 57.35 40.91 59.09 48.95 74.93 40.04 53.72 46.92 56.54 50.67 52.97
InkubaLM-0.4B 0 22.96 19.38 18.14 20.02 18.42 22.44 26.16 16.86 20.76 20.74 25.10 17.44 20.58 19.26 20.38 20.58
5 28.04 24.70 20.06 21.78 21.98 25.78 26.86 19.08 25.06 20.48 29.26 24.50 23.04 20.98 21.68 23.55
10 27.00 26.46 23.92 24.48 23.52 27.22 29.14 23.68 25.98 24.58 27.92 24.02 24.68 23.68 25.06 25.42
20 27.06 26.42 24.60 24.30 23.86 26.96 28.58 23.76 25.88 24.64 30.04 23.32 23.56 23.14 25.20 25.42
mt0-small 0 18.86 16.98 13.42 12.22 9.72 17.40 20.68 12.40 15.16 14.07 19.30 8.56 14.20 12.04 13.62 14.58
5 26.38 25.76 21.66 19.94 21.28 23.04 27.78 25.98 21.40 20.30 26.20 12.66 19.96 17.72 19.50 21.97
10 27.38 23.72 20.40 20.94 22.36 21.08 27.94 25.58 19.88 16.93 26.62 12.08 18.56 19.90 18.88 21.48
20 26.28 23.38 19.86 21.60 22.14 21.18 28.16 25.06 19.96 20.05 28.04 13.44 19.12 19.76 18.50 21.77
bloomz-7b1-mt 0 17.18 21.86 21.86 18.60 18.42 21.98 19.38 21.60 16.74 25.10 16.92 17.70 19.24 22.36 19.80 19.92
5 22.74 21.40 25.88 23.10 27.88 25.36 26.22 26.66 23.54 27.03 25.64 28.34 23.26 24.80 22.62 24.96
10 27.86 21.06 25.88 24.40 27.90 22.54 27.86 26.24 25.48 19.83 28.58 28.50 23.56 24.16 23.92 25.18
20 25.46 19.92 24.64 24.96 27.72 24.18 26.64 26.82 25.76 24.50 28.48 28.84 23.24 25.28 23.88 25.35
Mistral-7B-v0.1 0 21.18 17.46 11.98 6.84 8.04 15.18 20.82 6.76 8.16 10.64 25.62 8.38 9.72 8.22 8.92 12.53
5 37.68 31.60 44.54 34.90 36.84 46.98 39.46 50.00 34.12 49.58 35.64 32.04 34.24 41.04 32.54 38.75
10 39.68 36.50 50.56 37.82 36.94 45.72 39.02 52.10 31.82 54.34 35.84 30.82 32.64 42.90 32.46 39.94
20 36.64 35.46 52.94 38.44 38.98 52.28 39.24 54.60 33.28 55.98 37.42 30.44 34.12 46.40 35.00 41.41
aya-23-35B 0 17.04 19.20 17.20 20.34 14.04 21.34 20.38 20.02 18.56 21.20 17.10 10.68 20.32 17.48 20.24 18.34
5 37.78 56.40 57.00 39.70 42.70 48.78 39.24 57.90 38.72 56.32 35.68 40.24 36.76 48.88 44.58 45.38
10 38.50 53.96 61.82 44.42 45.80 52.82 42.16 59.26 37.60 61.08 35.88 39.92 36.18 48.44 45.00 46.86
20 38.10 52.90 63.76 46.14 46.52 57.62 42.26 61.68 38.68 73.68 38.12 38.54 37.74 51.52 44.52 48.79
Gemma-2-9B 0 27.42 28.36 25.46 14.48 15.00 24.08 23.74 24.74 10.92 32.12 25.52 19.50 13.40 21.66 12.52 21.26
5 56.60 57.68 61.14 49.98 48.12 60.26 45.10 60.30 46.42 72.88 48.40 46.50 35.82 55.96 46.60 52.78
10 57.84 61.08 61.72 56.62 53.78 63.54 45.78 60.54 51.72 76.50 45.48 48.16 38.82 60.22 52.58 55.63
20 60.96 59.94 64.38 55.90 54.56 64.14 44.70 62.54 52.22 77.56 44.94 48.92 39.02 60.50 53.68 56.26
Gemma-2-27B 0 41.78 41.24 41.12 28.96 31.36 42.08 33.46 49.78 31.10 54.54 30.84 27.74 25.64 41.46 28.02 36.61
5 59.62 64.14 65.62 54.62 56.14 61.08 46.76 61.78 54.12 77.96 52.12 47.26 40.48 61.72 54.28 57.18
10 60.70 61.36 64.88 58.90 56.84 64.86 46.96 61.60 52.86 80.04 49.82 50.86 41.90 59.66 56.94 57.88
20 62.28 59.86 65.80 59.60 58.76 66.06 48.90 63.24 52.90 81.18 53.36 50.34 43.08 59.10 56.88 58.76
Llama-3.1-8B 0 17.36 19.34 18.82 24.36 12.70 23.90 22.56 20.10 24.86 20.82 14.18 9.02 21.42 18.60 21.94 19.33
5 38.08 40.88 51.68 36.78 35.90 38.78 31.24 52.38 30.62 58.10 30.20 34.82 29.22 40.42 30.90 38.67
10 42.04 43.10 54.76 37.78 37.78 43.02 35.92 53.88 32.98 64.96 31.92 34.52 28.20 38.98 30.72 40.70
20 47.74 40.84 57.92 41.64 37.76 48.86 39.78 55.36 30.26 69.38 37.12 33.60 27.84 42.68 29.12 42.66
Llama-3.1-70B 0 36.34 43.34 42.64 35.64 32.52 38.52 31.54 48.66 31.14 49.36 27.08 25.54 28.98 40.64 35.50 36.50
5 58.52 66.24 62.88 52.86 52.88 58.14 45.50 64.38 52.72 73.48 44.42 43.90 39.02 58.08 55.24 55.22
10 61.18 64.46 62.20 56.40 55.10 59.00 46.60 63.58 54.40 76.50 49.62 45.76 39.04 57.96 55.74 56.50
20 60.38 61.36 63.80 57.40 53.80 62.48 49.02 63.18 51.82 77.72 53.90 47.50 40.34 57.94 56.36 57.13
GPT-4o 0 61.78 66.41 73.75 56.91 68.82 62.53 60.01 65.94 63.42 73.66 45.75 52.72 58.68 75.21 54.47 62.67
5 66.73 73.53 77.33 55.44 76.73 72.27 70.94 66.29 59.33 80.95 61.11 73.21 60.45 76.98 59.34 68.71
10 67.94 75.54 77.54 58.15 80.08 75.07 72.27 67.14 62.75 84.19 57.52 72.28 65.75 76.74 65.05 70.53
20 68.08 76.16 78.69 58.34 80.81 74.86 72.33 65.41 66.44 84.61 59.55 75.86 66.08 77.11 71.36 71.71
Lang. avg. - 40.80 41.73 44.47 37.60 39.26 43.51 37.59 44.99 36.16 51.01 36.37 35.05 33.03 41.56 36.43 39.97
Table 11: Model performance (Macro F1-score) for zero- and few shot classifiers across the 14 languages in AfriHate. Best performance for each language is highlighted in bold.
model #shots amh ary arq hau ibo kin oro pcm som swa tir twi xho yor zul Avg.
SetFit 0 33.20 42.92 26.63 24.02 38.61 21.29 28.72 39.23 29.26 35.26 47.32 52.15 37.62 41.27 35.03 35.50
5 45.11 35.91 31.89 50.81 50.30 51.26 36.50 52.54 49.93 54.83 31.63 38.40 46.14 31.62 35.71 42.84
10 49.26 37.05 44.27 50.43 40.80 48.60 42.56 56.87 40.27 62.88 39.74 49.28 42.60 49.21 42.03 46.39
20 50.33 40.92 53.25 54.34 65.65 57.28 43.08 58.44 48.86 74.68 38.95 55.87 46.78 53.85 51.10 52.89
InkubaLM-0.4B 0 39.74 23.62 30.26 22.66 23.04 30.96 41.14 19.84 28.30 27.02 47.24 23.16 27.30 23.70 27.38 29.02
5 39.12 23.72 31.88 25.28 34.36 32.60 44.88 23.74 30.86 25.62 48.12 43.96 28.14 26.88 27.52 32.45
10 37.20 29.52 32.64 38.38 27.74 37.58 43.26 28.82 41.26 31.22 45.84 32.90 34.22 31.54 35.76 35.19
20 37.64 29.18 32.68 38.64 28.70 37.84 43.06 28.98 41.34 31.37 48.88 32.30 33.82 30.84 36.12 35.43
mt0-small 0 38.14 23.36 31.76 28.70 17.02 35.56 43.74 22.70 31.34 27.47 42.80 14.66 28.32 23.16 27.62 29.09
5 36.70 24.90 34.12 29.18 29.48 31.36 41.10 32.06 25.26 27.90 42.56 16.20 33.04 28.38 32.86 31.01
10 37.40 25.74 35.06 26.36 33.44 28.30 40.46 31.98 24.00 23.37 43.04 15.88 31.32 29.72 32.00 30.54
20 36.50 25.98 35.08 26.20 34.50 28.44 40.30 31.54 23.86 23.40 44.76 17.08 31.68 28.92 32.30 30.70
bloomz-7b1-mt 0 34.62 34.38 33.08 32.66 29.18 36.20 37.30 32.76 33.90 38.53 35.82 28.88 31.94 33.70 32.24 33.68
5 32.36 37.34 46.74 37.52 62.56 30.88 31.96 48.14 33.90 40.97 36.68 68.50 39.36 48.86 39.64 42.36
10 35.16 36.96 45.92 37.72 62.40 27.74 33.06 47.10 35.04 36.97 38.70 67.94 38.56 47.66 39.44 42.02
20 33.16 35.60 44.18 37.88 63.12 29.12 31.82 47.04 34.78 37.93 38.72 68.84 38.14 49.08 39.08 41.90
Mistral-7B-v0.1 0 42.26 30.62 17.28 10.52 13.00 27.62 45.42 11.14 13.86 18.88 55.02 14.28 16.36 13.44 14.54 22.95
5 37.92 43.46 57.14 51.96 58.42 50.40 45.70 59.32 52.64 58.58 41.36 69.52 47.50 57.72 45.96 51.84
10 40.42 46.68 60.66 56.38 61.74 50.00 46.70 60.42 51.50 63.18 44.88 70.26 46.40 60.50 46.42 53.74
20 39.48 45.08 61.70 56.98 65.10 55.64 48.82 61.22 54.02 64.28 51.38 70.04 48.64 64.94 50.16 55.83
aya-23-35B 0 34.54 32.76 28.52 37.76 18.66 41.04 42.48 30.34 40.76 34.04 35.84 14.40 35.42 28.22 35.96 32.72
5 42.20 61.86 64.38 55.32 56.42 56.56 49.96 61.12 57.46 69.84 45.70 63.86 48.96 63.48 51.14 56.55
10 45.74 61.22 69.00 59.78 63.98 58.66 52.30 62.08 56.98 74.10 51.08 69.58 49.06 66.50 52.44 59.50
20 46.76 59.70 71.40 59.84 68.30 62.06 54.14 65.44 57.78 75.68 56.90 71.20 51.72 70.54 56.12 61.84
Llama-3.1-8B 0 29.14 34.10 36.04 55.82 22.60 48.60 42.96 39.84 55.00 49.04 23.18 15.20 46.18 38.02 47.42 38.88
5 41.50 50.12 63.90 46.56 64.48 42.24 34.94 61.22 39.08 63.88 35.80 68.12 43.66 58.14 43.30 50.46
10 45.16 51.70 66.04 47.72 66.12 44.60 39.16 62.70 36.90 68.54 39.88 69.84 43.20 60.88 42.68 52.34
20 50.20 50.56 67.92 48.18 66.92 48.50 44.10 64.28 38.70 72.00 48.26 70.40 42.72 62.24 43.16 54.54
Gemma-2-9B 0 48.26 37.78 26.88 17.60 16.24 34.34 46.86 26.68 16.22 37.12 57.98 27.20 18.20 22.96 16.58 30.06
5 61.10 63.62 65.72 55.20 62.84 63.24 56.34 63.36 54.46 74.78 62.12 65.44 46.24 70.14 52.12 61.11
10 61.26 63.78 65.98 60.64 66.38 64.60 56.40 64.08 57.28 78.16 61.20 69.76 49.66 73.02 54.74 63.13
20 64.14 61.86 67.90 60.26 69.38 65.48 57.24 64.50 58.62 79.24 62.98 70.68 50.02 73.96 56.00 64.15
Gemma-2-27B 0 53.40 46.38 44.10 33.62 31.08 49.02 50.82 54.20 38.16 59.88 55.58 31.34 29.66 45.96 31.90 43.67
5 61.40 66.00 69.40 60.44 68.30 63.26 56.34 64.44 62.34 80.18 59.90 64.34 51.36 75.60 56.24 63.97
10 62.44 64.76 69.74 64.70 70.62 66.22 56.24 64.22 61.24 81.94 59.52 70.82 53.18 75.34 59.16 65.34
20 64.12 63.60 70.24 65.74 72.84 67.72 59.18 66.06 63.20 82.98 64.44 73.28 56.04 76.50 61.18 67.14
Llama-3.1-70B 0 46.28 49.84 52.30 51.26 43.72 50.38 48.30 56.50 46.38 60.14 43.64 35.68 42.64 51.34 47.00 48.36
5 59.10 68.32 67.12 58.96 66.88 60.90 54.24 67.20 61.20 75.76 45.18 66.18 50.18 72.60 57.30 62.07
10 61.36 67.80 67.70 63.00 68.78 61.50 55.86 66.62 63.38 78.74 50.48 68.34 52.04 73.22 58.92 63.85
20 60.72 65.62 69.48 64.00 70.02 65.08 58.78 67.02 64.46 80.02 56.40 72.52 53.50 74.98 61.26 65.59
GPT-4o 0 61.70 56.05 66.45 45.90 44.21 44.92 44.93 49.67 38.50 57.24 46.01 33.90 36.72 48.62 36.34 47.41
5 66.69 73.26 71.03 53.24 74.23 72.39 65.30 55.93 56.19 79.02 57.64 61.36 51.81 69.65 58.45 64.41
10 67.09 75.12 73.81 55.28 78.02 74.47 66.97 56.47 58.34 72.83 56.50 61.11 61.61 70.78 65.12 66.23
20 67.22 76.05 75.00 56.08 78.17 74.85 66.01 50.33 61.95 84.14 56.91 51.67 62.78 72.22 61.84 66.35
Lang. avg. - 47.21 47.15 51.73 45.76 51.10 48.39 47.03 49.50 44.74 56.22 47.65 50.37 42.15 51.59 43.85 48.30
Table 12: Model Performances (Accuracy) for Zero and Few shot Learning
Figure 1: Annotation Guidelines and Definitions